# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v2

@@ III-A System Design
Aiming at the problem of dim object detection in aerial images under few-shot conditions, we propose a brain-eye-computer based object detection system. As shown in Fig. 2, the system is designed according to EEG, vision, and eye movement properties. Firstly, the computer processes the image data through the feature extractor and the region proposal network (RPN) to obtain the image with the pre-detection box (see Section III-B). Subsequently, this part of the image is presented in the display screen through the eye-tracking-based slow serial visual presentation (ESSVP) paradigm (see Section III-C), which in turn induces the subject to generate the corresponding EEG signals. Through the real-time eye movement data recorded by the eye tracker and the time-stamp synchronization signal of the trigger box, the computer simultaneously extracts the EEG data recorded by the EEG cap and the image of the subject’s attention area at the corresponding time. After that, the paired data will be preprocessed, and recognized by the proposed AMBOKD method (see Section IV) . Finally, according to the eye movement data, the position of the target image in the original image is found for target localization, so as to relize the target detection task.
The hardwares of the system primarily serves the roles of sensor data acquisition and algorithm execution, comprising a 30-inch display with a 2K resolution, a computer equipped with an NVIDIA GeForce GTX 2070s GPU, a 64-channel EEG cap, signal amplifier, trigger box, router, eye tracker, and various data cables. The display and computer are key components, utilized for the playback of ESSVP paradigm and the real-time storage and processing of EEG, eye movement, and image data. The displayer is positioned 50$\sim$70 cm directly in front of the subjects to ensure maximal elicitation of ERP signals and more accurate collection of eye movement data. Positioned directly below the display at a 15-degree angle towards the user’s eyes, the eye tracker effectively captures gaze positions, transmitting them to the computer at a frequency of 50Hz. EEG data is collected via 64 wet electrodes in the EEG cap, with the signals amplified by a connected amplifier before being transmitted to the computer. The trigger box synchronizes various types of event data (e.g., auditory, visual, and program outputs) with neurophysiological data with high event precision ($<1$ms), serving as a cornerstone for subsequent data analysis by synchronizing EEG data with event-triggered label signals.
Figure 3: The images with region proposals.
@@ III-B Suspicious Region Detection
The RPN is an integral component in deep learning and computer vision, particularly in object detection tasks [54]. Operating by processing a feature map, the RPN employs a sliding window approach across various locations, scales, and aspect ratios using an anchor boxes. This network evaluates the likelihood of each anchor enclosing an object and adjusts its position and dimensions accordingly, generating precise region proposals.
As illustrated in Fig. 2, the designed system uses a pretrained ResNet50 feature encoder to obtain feature maps of image data, and employs RPN to generate region proposals, which are denoted with black boxes. The anchors used have scales of 16, 32, 48, 64, and 80, with aspect ratios of 0.5, 0.75, 1.0, 1.5, and 2.0. We employ UAVs to capture images of toy models in open outdoor settings, forming the training set. Meanwhile, the test set comprises images taken in complex real-world environments. Fig. 3 shows the difference between the training set and the validation set. It can be seen that although the background of the validation set is complex and the target size is small, the RPN network trained by the training set can still effectively detect suspicious targets. These results confirm the effectiveness of the employed RPN network.
@@ III-C ESSVP Paradigm
To elicit ERPs more efficiently and obtain the EEG-image data, we propose the ESSVP paradigm according to the RSVP [28] and the AVEP [26] experimental paradigm. In this paradigm, visual stimuli are presented at a slow and controlled pace, providing the subjects sufficient time to search for specific targets. By integrating eye movement technology, we can accurately track the observer’s gaze and determine the specific areas of interest during the dim target recognition task.
Figure 4: Examples of stimulus materials.
The ESSVP paradigm firstly presents the experiment guidance for 1 min, and then the target example is shown to the subject for 3$\sim$4 s, as illustrated in Fig. 2. In this formal experiment, participants view 16 sequences of stimuli, each containing 50 images. These sequences are evenly divided into two sessions according to the region marking method. In each session, the first six sequences display images of toy models such as armored vehicles and airplanes, taken in simple scenes, with each image shown for 3 s. The last two sequences in each session present real images of real models, taken in complex scenes, with each image displayed for 4 s. The target in the first six sequences is the armored model, whereas in the last two sequences, the target is the vehicle. The different types of images used in the ESSVP paradigm are illustrated in Fig. 4.
The probability of the target image being included in each sequence is 40%. 1$\sim$2 nontarget images are randomly inserted between each target image to avoid the attentional blink caused by successive targets. During the stimulus presentation, we instruct subjects to actively search for the target among the candidate regions in the image, which is generated by the region proposal network in advance. Each candidate region must be fixated for 0.5 s. To prevent excessive visual fatigue due to prolonged task engagement, subjects are allowed to rest after each sequence, and the duration of the rest is determined by the subjects themselves.
@@
Figure 5: The experimental configuration of ESSVP.
The experimental environment for ESSVP is illustrated in Fig. 5. Specifically, the EEG signals sampled at a rate of 1000 Hz are collected through 64 wet electrodes adhering to the 10$\sim$20 standard [55] on the EEG cap. The impedance of these electrodes is maintained below 10k. Throughout the experiment, subjects are seated comfortably in a quiet environment, positioned approximately 60 cm away from the display screen, facing the displayed image. To ensure signal quality, and we instruct participants to maintain stability and minimize body and head movements while observing the visual presentation paradigm. This directive aims to minimize potential noise interference to the EEG signals.
@@ III-D Data Collection and Preprocessing
The ESSVP dataset used in this study includes EEG data from 10 subjects, all of whom are college students aged between 22 and 26, with normal vision and no history of mental illness. All subjects were provided with a clear understanding of the experimental procedure, task requirements, and brief information about the target characteristics before the experiment. In addition, the subjects signed an informed consent form. Ethical approval for the experiment was obtained from the relevant committee.
This dataset consists of 13405 samples, including 3880 positive samples with the target image and 9525 negative samples without the target image. Each sample includes EEG data from 59 electrode channels, as 5 out of the 64 electrodes are redundant and therefore excluded. The EEG data are filtered using a 2$\sim$30 Hz bandpass filter and downsampled from the original 1000 Hz to a rate of 250 Hz. The data is then segmented into 1.2 s (-500$\sim$700 ms) samples based on the trigger. Baseline correction is applied by employing the data from the first 200 ms interval as a reference to subtract the mean activity level. The image data used has a minimum pixel size of $150\times 150$, corresponding to the center of the focused candidate box. All images are resized to $224\times 224$ pixels to standardize the input for image network. The images in the validation set are preprocessed by introducing noise to simulate real-world conditions and evaluate the algorithm’s robustness, as shown in Fig. 6. Specifically, one-third of the images are enhenced with 0.2 Gaussian noise and another third of the images are enhenced with 0.2 salt and pepper noise.
@@ IV Methodology
The framework of the proposed adaptive modality balanced online knowledge distillation (AMBOKD) method is illustrated in Fig. 7. Initially, AMBOKD receives paired data $X$ and extract EEG and visual features through the visual encoder and EEG encoder. These features are subsequently passed through the multi-head self-attention module to generate the fusion features, and the features from each domain are sent to their respective classifiers. The features extracted from the visual, EEG, and fusion domains are classified and employed to compute the losses of the three modalities using the cross-entropy (CE) and KD loss functions. Subsequently, the three modalities take turns as students and teachers to compute the total loss, the weight of the teacher influence, and the dynamic student gradient for parameter updating. The processes of the feature extraction, feature fusion, the online knowledge distillation, and the adaptive modality balancing are detailed as follows.
@@ IV-A1 Visual Encoder
Three main components constitute EfficientNet-B0: convolutional neural networks, batch normalization functions, and mobile inverted bottleneck convolution, inspired by the MobileNet [19] concept. MBConv consists of three types of layers: convolutional layers, depth-wise convolutional layers, and squeeze-and-excitation networks (SENet) [56]. The convolutional layer expands and compresses channels, whereas the depth-wise convolutional layer reduces the parameter count. However, SENet focuses on channel relationships by assigning importance weights to various channels.
@@ IV-A2 EEG Encoder
The frequency encoder can effectively learn frequency-related features and capture the spectral characteristics of EEG signals by employing specific kernels in the multi-scale convolution module. Spatial representations are learned by the spatial encoder using a graph convolution module, which exploits the inherent spatial relationships between various electrode channels, enabling the spatial patterns present in the EEG data to be captured by MCGRAM. To extract temporal features and obtain the final global representation of the EEG modality, and the temporal encoder uses a two-layer long short-term memory network and a self-attention module. This combination allows temporal dependencies to be effectively modeled and crucial temporal patterns in the EEG signals to be captured by the model.
In summary, MCGRAM can capture and encode relevant information from EEG signals by cascading frequency, spatial, and temporal blocks. This resulted in a comprehensive and informative representation for further analysis and classification tasks. The representation of the EEG modality is computed as
@@
It should be noted that alternative algorithms can be employed as substitutes for the visual and EEG encoders. For example, algorithms such as ResNet-50 [18] and MobileNetV2 [19] can be used in the visual encoder, whereas EEGNet [57] and TSception [58] are viable options in the EEG encoder. The superiority and generalization of the AMBOKD approach proposed in this study can be effectively demonstrated by incorporating these alternative algorithms and conducting subsequent experimental validations (see Section V-E).
@@ IV-B Fusion of Modality Representation
Figure 8: Structure of the multi-head self-attention module for EEG and visual features fusion.
We develope a fusion model to integrate information from both EEG and visual modalities. This model is considered as novel modality that collaborates with the original modality in the mutual learning process. To focus on the most important information across various representations subspaces [59], a multihead self-attention module is employed within the fusion model. The creation of a global representation that captures the combined knowledge is facilitated by this module.
The multi-head self-attention module is designed to extract and align crucial information from EEG and visual modality features, as illustrated in Fig. 8. Dynamically learned weights are used to weigh and fuse the features of both modalities to obtain global fusion features.
Specifically, a fully connected (FC) layer is used to align the EEG feature $F_{e}\in\mathbb{R}^{f^{e}}$ and visual feature $F_{v}\in\mathbb{R}^{f^{v}}$ of each pair of data. To obtain the preliminary fused feature $F_{c}\in\mathbb{R}^{f^{c}}$, the aligned features are then concatenated. Three sets of FC layers were used to independently transform the features $F_{c}$ into queries $Q$, keys $K$, and values $V$ for the attention layer of the $j$-th head. The softmax function is used to multiply, scale, and normalize $Q$ and $K$ to learn the attention scores $A\in\mathbb{R}^{f^{c}\times f^{c}}$ between the two modalities:
@@
where ||||| | denotes the concat function and $J$ represents the number of heads.
The features obtained from the three domains are classified separately using their respective classifiers. For example, considering the EEG domain, the classification process can be described as follows:
$\text{G}_{e}=W_{e}{F_{e}}+b_{e},$ (6)
where $W_{e}$ denotes the weights and $b_{e}$ is the biases.
@@ IV-D2 Dynamic Ratios for Backward Gradients
$\phi_{v}(\theta_{v},F_{v}^{i})$ and $\phi_{e}(\theta_{e},F_{e}^{i})$ are simplified as $\phi_{v}^{i}$ and $\phi_{e}^{i}$ for convenience. When a particular modality, such as the visual modality, exhibits classification performance, it contributes more to the gradient $\frac{\partial\mathcal{L}_{CE}^{f}}{\partial G_{f}^{i}}$ through the expression $W_{f}W_{f}^{v}\phi_{v}^{i}$, thereby leading to a lower loss globally. Therefore, the EEG modality, characterized by lower confidence in accurate predictions, will obtain limited optimization during parameter updates via backpropagation. This phenomenon shows that the unimodal model might be overtrained or underoptimized when the training of the multimodal model is about to converge.
@@
EEGNet 70.65/6.55 69.12/6.07 71.68/4.98 69.84/6.14 $<10^{-3}$ 68.86/5.46 67.44/5.76 69.51/5.71 68.00/5.20 $<10^{-3}$
MCGRAM 82.47/7.11 75.18/8.47 76.69/6.55 76.04/9.48 $<10^{-3}$ 80.44/5.32 72.33/6.12 73.59/5.81 74.34/4.95 $<10^{-3}$
AMBOKD-E* 83.56/6.61 77.66/8.21 78.59/7.19 78.25/8.31 -- 81.85/5.71 75.95/5.75 76.71/5.09 77.05/5.07 --
ResNet-50 84.55/7.54 75.45/11.26 77.37/9.93 79.84/5.21 $<10^{-3}$ 84.42/6.63 76.04/7.43 77.04/7.36 79.53/6.03 $<10^{-3}$
EfficientNet 78.81/8.89 81.82/2.79 83.96/2.40 87.02/1.47 $<10^{-3}$ 84.13/3.92 80.68/4.63 82.81/4.15 86.53/2.34 $<10^{-3}$
AMBOKD-V* 88.54/3.50 83.08/2.52 84.88/2.25 87.54/1.49 -- 87.88/5.22 82.22/3.84 83.96/3.59 87.18/2.15 --
MLB 86.56/3.69 82.31/2.73 84.21/2.40 86.66/1.62 $<10^{-3}$ 84.74/4.42 80.95/7.85 82.74/6.13 85.20/3.64 $<10^{-3}$
MKD 90.21/4.45 82.87/2.45 84.73/2.21 87.54/1.39 $<10^{-3}$ 89.35/3.54 81.80/3.97 83.65/3.68 87.04/2.12 $<10^{-3}$
EMKD 90.51/4.22 82.65/2.42 84.57/2.18 87.42/1.34 $<10^{-3}$ 88.93/3.50 80.95/4.00 83.00/3.68 86.62/2.06 $<10^{-3}$
CA-MKD 90.52/4.19 82.78/2.52 84.67/2.24 87.46/1.42 $<10^{-3}$ 89.10/3.39 81.45/4.03 83.39/3.70 86.83/2.20 $<10^{-3}$
DML* 90.46/3.68 79.73/4.97 82.44/3.85 85.89/2.38 $<10^{-3}$ 88.87/3.21 77.66/5.50 80.60/4.52 84.84/2.87 $<10^{-3}$
KDCL* 89.91/4.22 79.05/4.72 81.74/3.71 84.55/2.25 $<10^{-3}$ 89.12/3.11 78.70/4.63 81.34/4.02 85.42/2.30 $<10^{-3}$
EML* 89.21/4.02 80.25/2.92 82.52/2.54 84.74/1.92 $<10^{-3}$ 90.64/3.06 80.89/4.06 82.93/3.74 86.40/2.16 $<10^{-3}$
AMBOKD (ours)* 93.66/2.95 83.33/2.51 85.08/2.24 87.82/1.38 -- 93.30/2.64 82.32/3.66 84.05/3.42 87.27/2.02 --

### v3

@@ III-A System Design
To address dim object detection in aerial images under few-shot conditions, we propose a brain-eye-computer based object detection system. As shown in Fig. 2, the system is designed according to EEG, vision, and eye movement properties. Firstly, the computer processes the image data through the feature extractor and the region proposal network (RPN) to obtain the image with the pre-detection box (see Section III-B). Subsequently, this part of the image is presented on the display screen through the eye-tracking-based slow serial visual presentation (ESSVP) paradigm (see Section III-C), which in turn induces the subject to generate the corresponding EEG signals. Through the real-time eye movement data recorded by the eye tracker and the time-stamp synchronization signal of the trigger box, the computer simultaneously extracts the EEG data recorded by the EEG cap and the image of the subject’s attention area at the corresponding time. The paired data is then preprocessed and analyzed using the proposed AMBOKD method (see Section IV). Finally, using eye movement data, the target’s position in the original image is identified to achieve target detection.
Figure 3: Examples of stimulus materials.
@@ III-B Suspicious Region Detection
The RPN is an integral component in deep learning and computer vision, particularly in object detection tasks [37]. Operating by processing a feature map, the RPN employs a sliding window approach across various locations, scales, and aspect ratios using an anchor box. This network evaluates the likelihood of each anchor enclosing an object and adjusts its position and dimensions accordingly, generating precise region proposals. The designed system uses a pre-trained ResNet-50 feature encoder [16] to obtain feature maps of image data, and employs RPN to generate region proposals, which are denoted with black boxes. The anchors used have scales of 16, 32, 48, 64, and 80, with aspect ratios of 0.5, 0.75, 1.0, 1.5, and 2.0.
@@ III-C ESSVP Paradigm
To elicit ERPs more efficiently and obtain the EEG-image data, we propose the ESSVP paradigm according to the RSVP [21] and the AVEP [20] experimental paradigm. Compared to the RSVP paradigm, the ESSVP paradigm offers improved processing of complex UAV-captured imagery, enabling detection of dim targets rather than conspicuous ones. In this paradigm, visual stimuli are presented at a slow and controlled pace, providing the subjects sufficient time to search for specific targets. By integrating eye movement technology, we can accurately track the observer’s gaze and determine the specific areas of interest during the dim target recognition task.
Figure 4: The experimental configuration of ESSVP.
The ESSVP paradigm first presents experimental guidance for 1 minute, followed by displaying the target example to the subject for 3–4 seconds, as shown in Fig. 2. In this formal experiment, participants view 16 sequences of stimuli, each containing 50 images. These sequences are evenly divided into two sessions according to the region marking method. In each session, the first six sequences display images of toy models such as armored vehicles and airplanes, taken in simple scenes, with each image shown for 3 s. The last two sequences in each session present real images of real models, taken in complex scenes, with each image displayed for 4 s. The target in the first six sequences is the armored model, whereas in the last two sequences, the target is the vehicle. The different types of images used in the ESSVP paradigm are illustrated in Fig. 3.
[…]
@@
[…]
@@ III-D Data Collection and Preprocessing
The ESSVP dataset includes EEG data from 10 subjects (6 males, 4 females) recruited through stratified sampling to ensure demographic diversity. All subjects were right-handed college students aged 22-26 (mean=23.4, SD=1.2) with normal or corrected-to-normal vision and no history of mental illness. All subjects were naive to EEG paradigms and were provided with a clear understanding of the experimental procedure, task requirements, and brief information about the target characteristics before the experiment. In addition, the subjects signed an informed consent form. Ethical approval for the experiment was obtained from the relevant committee.
[…]
@@ IV Methodology
The framework of the proposed adaptive modality balanced online knowledge distillation (AMBOKD) method is illustrated in Fig. 6. In this method, we use two feature extractors and a multi-head self-attention mechanism (see Sections IV-A and IV-B) to process EEG and visual data. Then, we introduce the mutual learning with online knowledge distillation and adaptive modality balancing strategies to optimize the fusion model for better performance. (Section IV-C and Section IV-D)
Initially, AMBOKD receives paired data $X$ and extracts EEG and visual features through the visual encoder and EEG encoder. These features are subsequently passed through the multi-head self-attention module to generate the fusion features, and the features from each domain are sent to their respective classifiers. The features extracted from the visual, EEG, and fusion domains are classified and employed to compute the losses of the three modalities using the cross-entropy (CE) and KD loss functions. Subsequently, the three modalities take turns as students and teachers to compute the total loss, the weight of the teacher influence, and the dynamic student gradient for parameter updating. The processes of the feature extraction, feature fusion, the online knowledge distillation, and the adaptive modality balancing are detailed as follows.
@@ IV-A2 EEG Encoder
The frequency encoder can effectively learn frequency-related features and capture the spectral characteristics of EEG signals by employing specific kernels in the multi-scale convolution module. Spatial representations are learned by the spatial encoder using a graph convolution module, which exploits the inherent spatial relationships between various electrode channels. To extract temporal features and obtain the final global representation of the EEG modality, the temporal encoder uses a two-layer long short-term memory network and a self-attention module to model the temporal dependencies and crucial temporal patterns.
In summary, MCGRAM captures and encodes relevant information from EEG signals by cascading frequency, spatial, and temporal blocks. This resulted in a comprehensive and informative representation for further analysis and classification tasks. The representation of the EEG modality is computed as
@@ IV-B Fusion of Modality Representation
We develop a fusion model to integrate EEG and visual modality information. This model is considered as a novel modality that collaborates with the original modality in the mutual learning process. To focus on the most important information across various representations subspaces [40], a multihead self-attention module is employed within the fusion model. This module creates a global representation capturing the combined knowledge.
The multi-head self-attention module is designed to extract and align crucial information from EEG and visual modality features. Dynamically learned weights are used to weigh and fuse the features of both modalities to obtain global fusion features. Specifically, a fully connected (FC) layer is used to align the EEG feature $F_{e}\in\mathbb{R}^{f^{e}}$ and visual feature $F_{v}\in\mathbb{R}^{f^{v}}$ of each pair of data. To obtain the preliminary fused feature $F_{c}\in\mathbb{R}^{f^{c}}$, the aligned features are then concatenated. Three sets of FC layers were used to independently transform the features $F_{c}$ into queries $Q$, keys $K$, and values $V$ for the attention layer of the $j$-th head. The softmax function is used to multiply, scale, and normalize $Q$ and $K$ to learn the attention scores $A\in\mathbb{R}^{f^{c}\times f^{c}}$ between the two modalities:
@@
where $||$ denotes the concat function and $J$ represents the number of heads.
The features obtained from the three domains are classified separately using their respective classifiers. For example, considering the fusion domain, the classification process can be described as follows:
$\text{G}_{f}=W_{f}{F_{c}}+b_{f},$ (6)
where $W_{f}$ denotes the weights and $b_{f}$ is the biases.
@@ IV-D2 Dynamic Ratios for Backward Gradients
$\phi_{v}(\theta_{v},F_{v}^{i})$ and $\phi_{e}(\theta_{e},F_{e}^{i})$ are simplified as $\phi_{v}^{i}$ and $\phi_{e}^{i}$ for convenience. When a particular modality, such as the visual modality, exhibits classification performance, it contributes more to the gradient $\frac{\partial\mathcal{L}_{CE}^{f}}{\partial G_{f}^{i}}$ through the expression $W_{f}W_{f}^{v}\phi_{v}^{i}$, thereby leading to a lower loss globally. Therefore, the EEG modality, characterized by lower confidence in accurate predictions, will obtain limited optimization during parameter updates via backpropagation. This phenomenon indicates that unimodal models may become overtrained or underoptimized as the multimodal model approaches convergence.
@@
EEGNet 70.65/6.55 69.12/6.07 71.68/4.98 69.84/6.14 $10^{-3}$ 68.86/5.46 67.44/5.76 69.51/5.71 68.00/5.20 $10^{-3}$
MCGRAM 82.47/7.11 75.18/8.47 76.69/6.55 76.04/9.48 $10^{-3}$ 80.44/5.32 72.33/6.12 73.59/5.81 74.34/4.95 $10^{-3}$
AMBOKD-E* 83.56/6.61 77.66/8.21 78.59/7.19 78.25/8.31 - 81.85/5.71 75.95/5.75 76.71/5.09 77.05/5.07 -
ResNet-50 84.55/7.54 75.45/11.26 77.37/9.93 79.84/5.21 $10^{-3}$ 84.42/6.63 76.04/7.43 77.04/7.36 79.53/6.03 $10^{-3}$
EfficientNet 78.81/8.89 81.82/2.79 83.96/2.40 87.02/1.47 $10^{-3}$ 84.13/3.92 80.68/4.63 82.81/4.15 86.53/2.34 $10^{-3}$
AMBOKD-V* 88.54/3.50 83.08/2.52 84.88/2.25 87.54/1.49 - 87.88/5.22 82.22/3.84 83.96/3.59 87.18/2.15 -
MKD 90.21/4.45 82.87/2.45 84.73/2.21 87.54/1.39 $10^{-3}$ 89.35/3.54 81.80/3.97 83.65/3.68 87.04/2.12 $10^{-3}$
EMKD 90.51/4.22 82.65/2.42 84.57/2.18 87.42/1.34 $10^{-3}$ 88.93/3.50 80.95/4.00 83.00/3.68 86.62/2.06 $10^{-3}$
CA-MKD 90.52/4.19 82.78/2.52 84.67/2.24 87.46/1.42 $10^{-3}$ 89.10/3.39 81.45/4.03 83.39/3.70 86.83/2.20 $10^{-3}$
C2KD 87.75/4.13 82.33/2.82 84.34/2.46 87.28/1.52 $10^{-3}$ 86.63/4.35 82.08/4.05 83.87/3.72 87.18/2.15 $10^{-3}$
DIST+ 90.39/4.19 83.42/5.21 84.57/5.15 86.93/2.78 $10^{-3}$ 89.77/3.66 83.12/4.37 84.43/4.27 86.92/2.83 $10^{-3}$
DML* 90.46/3.68 79.73/4.97 82.44/3.85 85.89/2.38 $10^{-3}$ 88.87/3.21 77.66/5.50 80.60/4.52 84.84/2.87 $10^{-3}$
KDCL* 89.91/4.22 79.05/4.72 81.74/3.71 84.55/2.25 $10^{-3}$ 89.12/3.11 78.70/4.63 81.34/4.02 85.42/2.30 $10^{-3}$
EML* 89.21/4.02 80.25/2.92 82.52/2.54 84.74/1.92 $10^{-3}$ 90.64/3.06 80.89/4.06 82.93/3.74 86.40/2.16 $10^{-3}$
MLEMKD* 89.12/2.02 80.32/2.28 82.83/2.14 86.28/1.24 $10^{-3}$ 88.45/2.04 79.31/4.28 81.76/3.94 85.85/2.12 $10^{-3}$
AMBOKD (ours)* 93.66/2.95 83.33/2.51 85.08/2.24 87.82/1.38 - 93.30/2.64 82.32/3.66 84.05/3.42 87.27/2.02 -